OpenStudy (anonymous):

will fan and medal Find x if f(x) = 2x + 7 and f(x) = -1. -4 -3 5

OpenStudy (anonymous):

F(x) = 2x +7 = -1 2x = -1 -7 2x = -8 X = -8/2 X = -4